knowing following information for a sample of 16: m = 8; Xbar = 10.22; sigma x...

  1. knowing following information for a sample of 16: m = 8; Xbar = 10.22; sigma x = 1.8. What is the z-score for this sample mean?

  2. a.

    1.23

    b.

    -1.23

    c.

    4.93

    d.

    -4.93

    e.

    None of the above
Answer #1

Solution :

Given ,

mean = = 8

standard deviation = = 1.8

n=16

=10.22

= 8

= / n = 1.8/16=0.45

using z-score formula

z = - /

z=10.22-8/0.45

z=4.93